OpenStudy (anonymous):

If volume of a sphere increases by 72.8% , how much percent does it's Surface Area increases ?? If volume of a sphere increases by 72.8% , how much percent does it's Surface Area increases ?? @Mathematics

OpenStudy (anonymous):

You want to use scaling ratios (k). You know the volume is 72.8% bigger, or 1.728 times bigger. This means that k^3=1.728 (because volume is k^3, area is k^2 and length is just k). If k^3=1.728, then k=1.2 and k^2=1.44. So the surface area is 1.44 times more, or 144%